OpenStudy (anonymous):

1/sqrt(2 + sqrt(3)) + 1/sqrt(3 + sqrt(4)) + .... + 1/sqrt(1000000 + sqrt(1000001)) = sqrt( a - sqrt(b)). a + b = .... ?

OpenStudy (anonymous):

@ganeshie8 and @dan815 help me ... how to simplify this summation of radicals :D

OpenStudy (inkyvoyd):

rationalize the denominator... some stuff should cancel out :)

OpenStudy (kainui):

Totally do what inky says, I just thought I'd share some stuff anyways, kinda incomplete but if you're curious I can explain more. My systematic way of solving these lately has been to write it in summation notation: \[\sum_{n=a}^b g(n)\] then look for solutions to: \[g(n) = f(n+1)-f(n)\] or \[g(n) = f(n)-f(n-1)\] The point is, then your sum becomes a telescoping series so you only have to evaluate the end points. Really when written out, it's basically the fundamental theorem of calculus: \[\sum_{n=a}^{b-1} \Delta f(n) = f(b)-f(a)\]

ganeshie8 (ganeshie8):

rationalizing doesn't help much here

OpenStudy (wcrmelissa2001):

not sure how to solve this but I think normally how it works for these kind of questions is that you simplify/rationalise/something the first one or two numbers and look for a pattern. This isn't very helpful sorry but I did a similar question before and found that each number actually was 1......yea....

OpenStudy (kainui):

@tanjung Are you sure you wrote the question correctly?

OpenStudy (anonymous):

Yes. That is right equations...

OpenStudy (anonymous):

Sorry i cant write it in latex. I want show you in picture, but i cant do it by using android

OpenStudy (kainui):

\[\sum_{n=2}^{10^6} \frac{1}{\sqrt{n+\sqrt{n+1}}} = \sqrt{a - \sqrt{b}}\] What is \(a+b\)

OpenStudy (anonymous):

Yep. That is like above. Thanks :D

OpenStudy (kainui):

How'd you get this question this seems tough lol

OpenStudy (anonymous):

My friend 's question.. i tried help her, but stuck also :v